OpenStudy (anonymous):

An isosceles trapezoid has bases of lengths 9 and 15 and a height of 10. Find the area of the trapezoid

OpenStudy (anonymous):

someone help me

OpenStudy (shamim):

Can u draw ur trapezoud???

OpenStudy (shamim):

*trapezoid

OpenStudy (anonymous):

it just tellme to find the area of trapezoid

OpenStudy (shamim):

Area=.5*(sum of 2 parallel sides)*height

OpenStudy (anonymous):

so it 5

OpenStudy (shamim):

Given Sum of 2 parallel sides=9+15=?

OpenStudy (shamim):

Height=10m

OpenStudy (anonymous):

24

OpenStudy (shamim):

No

OpenStudy (anonymous):

I add 9+15=24

OpenStudy (shamim):

Area of ur trapezoid=.5*(9+15)*10=.5*24*10=12*10=?

OpenStudy (anonymous):

than

OpenStudy (butterflydreamer):

 the formula for a trapezoid is: